Ling invested some money at 5% interest. Ling also invested $196 more than 3 times that amount at 11%. How much is invested at each rate if Ling receives $1005.76 in interest after one year? (Round to two decimal places if necessary.)
Answer by ankor@dixie-net.com(22740)

Ling invested some money at 5% interest.
Ling also invested $196 more than 3 times that amount at 11%.
How much is invested at each rate if Ling receives $1005.76 in interest after one year? (Round to two decimal places if necessary.)
:
let x = amt invested at 5%
" Ling also invested $196 more than 3 times that amount at 11%." therefore
(3x+196) = amt invested at 11%
:
.05x + .11(3x+196) = 1005.76
.05x + .33x + 21.96 = 1005.76
.38x = 1005.76 - 21.96
x = 984.20/.38
x = $2590 invested at 5%
then
3(2590) + 196 = $7966 invested at 11%
:
:
Confirm this solution
.05(2590) + .11(7966) = 1005.76